# SpatialBench: Benchmarking Multimodal Large Language Models for Spatial Cognition

### v1

@@ 4.2 Benchmark Construction
Qwen3-VL-235B-A22B-Instruct 38.84 37.79 2 43.76 70.10 40.43 34.33 68.37 24.49 30.30 32.99 31.00 36.08 25.51 37.50 33.33 26.67 43.30
48.68 37.40 32.99 33.33 41.07
Qwen3-VL-235B-A22B-Thinking 36.60 36.27 3 45.36 71.03 43.18 27.73 45.92 31.58 31.63 20.41 37.00 44.33 17.35 27.08 28.28 40.00 45.36
47.70 33.33 29.55 28.28 44.64
Qwen2.5-VL-72B-Instruct 32.73 33.82 4 26.14 57.94 43.48 30.21 37.76 13.27 27.27 36.73 30.00 40.21 16.33 43.75 36.36 6.67 37.11
38.33 29.01 33.33 36.36 33.04
Qwen3-235B-A22B-Instruct-2507 34.88 33.65 5 37.03 54.85 46.52 29.48 74.49 63.27 24.24 23.47 19.00 35.05 7.14 18.75 33.33 0.00 41.24
40.85 40.77 20.27 33.33 35.71
Qwen3-235B-A22B-Thinking-2507 34.99 33.25 6 32.00 60.90 24.29 32.13 58.24 28.26 29.67 39.13 23.91 28.09 40.91 21.84 29.55 40.00 33.70
40.24 35.81 30.30 29.55 34.58
Qwen3-VL-30B-A3B-Instruct 37.07 32.83 7 48.02 59.28 63.04 34.85 74.49 58.16 30.30 25.51 21.00 35.05 6.12 22.92 29.29 0.00 38.14
48.52 41.78 21.31 29.29 33.04
Qwen3-VL-30B-A3B-Thinking 36.83 31.22 8 48.79 63.40 27.83 30.72 71.43 45.36 31.31 27.55 32.00 35.05 11.22 32.29 26.26 13.33 30.93
46.20 41.46 26.12 26.26 28.57
Qwen3-Next-80B-A3B-Instruct 32.48 30.89 10 30.20 48.35 22.17 25.77 73.47 9.18 34.34 30.61 23.00 50.52 12.24 30.21 25.25 13.33 36.08
33.81 34.08 30.93 25.25 33.04
GLM-4V-Plus-0111 36.89 30.40 11 45.70 65.98 40.45 39.90 60.20 55.67 27.84 36.73 24.24 34.02 14.29 22.92 36.73 13.33 18.56
49.81 40.90 23.71 36.73 17.86
ERNIE-4.5-Turbo-VL-32K 34.72 28.05 13 44.06 64.85 24.78 28.45 72.45 60.20 25.25 27.55 23.00 24.74 10.20 27.08 30.30 0.00 21.65
44.25 41.58 20.62 30.30 18.75
GLM-4.5V 35.34 27.72 14 48.53 66.60 23.04 41.75 74.49 28.87 32.32 29.59 26.00 33.68 8.16 33.68 28.12 26.67 15.46
50.16 38.21 25.00 28.12 16.96
MiniMax-Text-01 33.49 27.59 15 50.79 61.65 32.61 35.57 45.92 40.82 24.24 44.90 24.00 39.18 3.06 25.00 28.28 6.67 21.65
48.14 35.90 22.34 28.28 19.64
QwQ-32B 29.76 23.81 16 37.82 54.54 30.87 24.33 72.45 35.71 17.17 31.63 18.00 34.02 9.18 11.46 28.28 13.33 13.40
38.30 34.89 18.21 28.28 13.39
Qwen2.5-VL-7B-Instruct 17.84 17.45 18 19.21 28.04 17.39 12.37 18.37 11.34 11.11 16.33 20.20 15.05 14.74 23.96 18.95 33.33 13.54
19.69 15.48 17.96 18.95 16.22
Table 1: Evaluation results on SpatialBench.
Obj.count
Obj.Size
Room Size
Abs. Distance
App. Order
App. Order (Self-Def Route)
Rel. Distance
Rel. Direction
Rel. Count
Multi-Hop Reasoning
Affordance
Landmark Constrained Location
Causal Reasoning
Visual Based Commands
Route Plan
Model Ave. Score Observation Topology & Relation Symbolic Reasoning Causality Planning
Gemini-2.5-pro 71.86 75.79 57.13 72.78 48.26 38.66 69.39 62.24 78.79 62.24 91.00 73.20 94.90 86.46 80.81 73.33 74.23
55.63 72.82 84.88 80.81 74.11
Gemini-2.5-pro (one-shot) 67.43 68.12 53.10 76.29 43.91 42.06 52.04 54.08 74.75 61.22 91.00 66.67 91.84 81.25 66.57 66.67 67.01
56.15 66.73 80.00 66.67 66.96
GPT-5-chat-latest 23.93 22.45 38.33 72.15 42.61 36.48 5.32 3.19 29.47 17.02 19.79 16.13 20.43 10.87 20.00 13.33 25.81
48.48 15.01 15.83 20.00 24.07
GPT-5-chat-latest (one-shot) 59.03 61.08 40.10 71.34 53.04 38.66 58.16 48.98 58.59 39.80 74.00 53.12 85.71 78.12 56.57 53.33 65.98
50.16 55.98 72.41 56.57 64.29
Qwen3-VL-235B-A22B-Instruct 38.84 37.79 43.76 70.10 40.43 34.33 68.37 24.49 30.30 32.99 31.00 36.08 25.51 37.50 33.33 26.67 43.30
48.68 37.40 32.99 33.33 41.07
Qwen3-VL-235B-A22B-Instruct (one-shot) 61.26 63.20 44.80 69.38 46.52 27.94 65.31 43.88 74.75 59.18 73.00 52.08 86.73 79.17 53.54 53.33 73.20
47.29 63.29 72.76 53.54 70.54
Table 2: Evaluation results under the one-shot setting.
Obj.count
Obj.Size
Room Size
Abs. Distance
App. Order
App. Order (Self-Def Route)
Rel. Distance
Rel. Direction
Rel. Count
Multi-Hop Reasoning
Affordance
Landmark Constrained Location
Causal Reasoning
Visual Based Commands
Route Plan
Model Ave. Score Observation Topology & Relation Symbolic Reasoning Causality Planning
Human Level 93.89 96.40 100.00 53.75 67.50 60.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00
Gemini-2.5-pro 71.86 75.79 57.13 72.78 48.26 38.66 69.39 62.24 78.79 62.24 91.00 73.20 94.90 86.46 80.81 73.33 74.23
Table 3: Human benchmark test results.
@@ 5.1 Setup
Models. We evaluate a diverse set of MLLMs to comprehensively assess their spatial cognitive abilities. Our benchmark covers both proprietary and open-source models with support for video understanding. For proprietary models, we include leading systems such as Gemini [gemini2023gemini], GPT [hurst2024gpt], and Claude-Sonnet. For open-source models, we evaluate representative families including Qwen [bai2023qwen], GLM [glm2024chatglm], MiniMax [li2025minimax], and ERNIE [zhang2019ernie], encompassing variants across a broad range of architectures, parameter sizes (7B–235B), and training paradigms, enabling a systematic comparison of how different modeling strategies influence spatial cognition performance in SpatialBench.
Metrics. Following [yang2025thinking], we adopt evaluation metrics tailored to the answer type. Specifically, tasks in SpatialBench are categorized into Multiple-Choice Answer (MCA) and Numerical Answer (NA) formats. For MCA tasks, we employ accuracy (ACC) as the primary evaluation metric [fu2025videomme, hendrycks2021measuring], which measures the proportion of exactly matched answers between the model’s predictions and the ground-truth labels. While for NA tasks, where answers involve continuous numerical values, we adopt the mean relative accuracy (MRA) metric [yang2025thinking, lin2014microsoft], defined as:
@@ 5.2 Benchmarking MLLM Performance
Table 1 presents the model performance on SpatialBench across five hierarchical levels. Proprietary models (e.g., Gemini-2.5-pro) occupy the top of the ranking and show substantially higher overall scores than most open-source counterparts. This gap is most pronounced on high-level tasks such as symbolic reasoning, causality, and planning. Nevertheless, several open-source series (e.g., Qwen variants) reach competitive performance on lower and mid levels, indicating that open models can achieve strong perceptual and relational understanding with carefully designed training paradigms. For open-source models, a clear correlation appears between model scale and average performance: larger models generally achieve higher overall scores, showing that scale remains an important factor. However, size alone does not guarantee better performance. Within the same model family, different versions (such as instruction-tuned vs. thinking mode) can show clear performance gaps. This suggests that architectural design and the integration strategy for visual and linguistic information plays a crucial role in spatial cognitive ability.
Figure 5 illustrates the average performance of several representative models across the five cognitive levels. It can be observed that a cross the board, observation and topology tasks are relatively easier for MLLMs. Most models achieve substantially higher scores on tasks such as object counting, size and distance estimation, and simple topological queries. By contrast, higher-level abilities remain challenging for models, including symbolic reasoning, causality, and visual planning. On these tasks, average performance drops noticeably and variance between models increases. This pattern suggests that while current MLLMs can reliably extract visual evidence and reason about basic relations, they struggle to (a) convert perceptual inputs into robust symbolic rules, (b) infer causal or dynamic consequences accurately, and (c) generate multi-step and convincing plans for a given objective.
@@ 5.3 One-Shot Evaluation
We select several representative models from the one-shot evaluation, including two proprietary models: Gemini-2.5-pro (the best-performing) and GPT-5-chat-latest (the weakest), as well as a strong open-source model Qwen3-VL-235B-A22B-Instruct. Gemini-2.5-pro shows a decline in performance, primarily across the intermediate cognitive levels. In contrast, GPT-5-chat-latest and Qwen3-VL demonstrate substantial improvements. Although these models still trail Gemini in absolute scores, their performance under prompting approaches that of the previously strongest model. This result indicates that even lower-performing models can rapidly enhance specific spatial reasoning abilities when given minimal in-context guidance, highlighting the potential of one-shot prompting to boost MLLMs’ higher-level cognitive capabilities. The results suggest that some powerful models, such as Gemini 2.5 Pro, can reason efficiently and accurately from context on their own, performing better than when given human-guided prompts. This indicates that their built-in contextual understanding is particularly well-suited for spatial intelligence tasks. By contrast, GPT‑5 appears comparatively weaker in intrinsic spatial logic, yet achieves more advanced spatial intelligence by leveraging the GPT series’ steadily improving linguistic capabilities [achiam2023gpt]. Similarly, Qwen3-VL, which demonstrates strong text-based reasoning skills [qwen3vl-hf], benefits from one-shot prompting in a way comparable to GPT‑5, highlighting the role of linguistic reasoning in enhancing spatial cognition under minimal guidance.
@@ 5.4 Benchmarking Human Performance
To quantify the gap between current MLLMs and human-level intelligence across different cognitive dimensions, we conduct a human benchmark experiment. In this experiment, 33 human participants are presented with SpatialBench, and they provide answers directly without guidance, allowing us to measure their unaided spatial reasoning and planning abilities. The results shown in Table 3 indicate that human achieve near-perfect performance across nearly all tasks with an overall score of 96.40. Humans perform particularly well on higher-level tasks: symbolic reasoning, causality, and planning, and all achieve essentially 100% accuracy, while even lower-level observation and topology tasks maintain strong performance. Across all levels, humans exhibit not only higher absolute accuracy but also more consistent performance, reflecting robust generalization and contextual understanding that MLLMs have yet to fully achieve. These findings underscore the substantial gap that still exists between human and machine spatial intelligence. While MLLMs show promising abilities in extracting visual information and reasoning about simple topological relations, they remain far from matching human performance in tasks that require integrating high-level reasoning. This test thus provides a clear target for future model development, highlighting the need to enhance multi-step reasoning and context-sensitive planning in MLLMs.
Figure 6 shows an example of the differences between Gemini and human. The human focuses on the key directional cue: the turning path of the white Volvo, and quickly eliminates irrelevant options based on spatial orientation, demonstrating strong goal-directed and spatially grounded reasoning. In contrast, Gemini describes the entire scene in a more exhaustive but unfocused manner, mentioning many vehicles and areas without identifying the crucial spatial relationship. This suggests that while MLLMs can recognize objects and describe scenes accurately, they often lack selective attention and directional understanding, leading them to infer by association rather than by reasoning about movement and geometry.

### v3

@@ 4.2 Benchmark Construction
Qwen3-VL-235B-A22B-Instruct 43.76 70.10 40.43 34.33 68.37 24.49 30.30 32.99 31.00 36.08 25.51 37.50 33.33 26.67 43.30
Qwen3-VL-235B-A22B-Thinking 45.36 71.03 43.18 27.73 45.92 31.58 31.63 20.41 37.00 44.33 17.35 27.08 28.28 40.00 45.36
Qwen2.5-VL-72B-Instruct 26.14 57.94 43.48 30.21 37.76 13.27 27.27 36.73 30.00 40.21 16.33 43.75 36.36 6.67 37.11
Qwen3-VL-30B-A3B-Instruct 48.02 59.28 63.04 34.85 74.49 58.16 30.30 25.51 21.00 35.05 6.12 22.92 29.29 0.00 38.14
Qwen3-VL-30B-A3B-Thinking 48.79 63.40 27.83 30.72 71.43 45.36 31.31 27.55 32.00 35.05 11.22 32.29 26.26 13.33 30.93
GLM-4V-Plus-0111 45.70 65.98 40.45 39.90 60.20 55.67 27.84 36.73 24.24 34.02 14.29 22.92 36.73 13.33 18.56
ERNIE-4.5-Turbo-VL-32K 44.06 64.85 24.78 28.45 72.45 60.20 25.25 27.55 23.00 24.74 10.20 27.08 30.30 0.00 21.65
GLM-4.5V 48.53 66.60 23.04 41.75 74.49 28.87 32.32 29.59 26.00 33.68 8.16 33.68 28.12 26.67 15.46
MiniMax-Text-01 50.79 61.65 32.61 35.57 45.92 40.82 24.24 44.90 24.00 39.18 3.06 25.00 28.28 6.67 21.65
QwQ-32B 37.82 54.54 30.87 24.33 72.45 35.71 17.17 31.63 18.00 34.02 9.18 11.46 28.28 13.33 13.40
Qwen2.5-VL-7B-Instruct 19.21 28.04 17.39 12.37 18.37 11.34 11.11 16.33 20.20 15.05 14.74 23.96 18.95 33.33 13.54
Table 3: Detailed evaluation results on SpatialBench for 15 tasks.
Model & Setting Score Ave. Obs. Topo. Symb. Caus. Plan.
Gemini-2.5-pro
Default 75.79 71.86 55.63 72.82 84.88 80.81 74.11
Few-shot 68.12 (-7.7) 67.43 (-4.4) 56.15 (+0.5) 66.73 (-6.1) 80.00 (-4.9) 66.67 (-14.1) 66.96 (-7.2)
GPT-5-chat-latest
Default 22.45 23.93 48.48 15.01 15.83 20.00 24.07
Few-shot 61.08 (+38.6) 59.03 (+35.1) 50.16 (+1.7) 55.98 (+41.0) 72.41 (+56.6) 56.57 (+36.6) 64.29 (+40.2)
Qwen3-VL-Inst
Default 37.79 38.84 48.68 37.40 32.99 33.33 41.07
Few-shot 63.20 (+25.4) 61.26 (+22.4) 47.29 (-1.4) 63.29 (+25.9) 72.76 (+39.8) 53.54 (+20.2) 70.54 (+29.5)
Table 4: Evaluation result on few-shot setting with $\Delta$ change. Bold indicates positive delta over 10 points on few-shot setting.
@@ 5.1 Setup
Models. We evaluate a diverse set of MLLMs to comprehensively assess their spatial cognitive abilities. Our benchmark covers both proprietary and open-source models with support for video understanding. For proprietary models, we include leading systems such as Gemini [56], GPT [30], and Claude-Sonnet. For open-source models, we evaluate representative families including Qwen [4], GLM [25], MiniMax [33], and ERNIE [74], encompassing variants across a broad range of architectures, parameter sizes (7B–235B), and training paradigms, enabling a systematic comparison of how different modeling strategies influence spatial cognition performance in SpatialBench.
Metrics. Following [66], we adopt evaluation metrics tailored to the answer type. Specifically, tasks in SpatialBench are categorized into Multiple-Choice Answer (MCA) and Numerical Answer (NA) formats. For MCA tasks, we employ accuracy (ACC) as the primary evaluation metric [22, 28], which measures the proportion of exactly matched answers between the model’s predictions and the ground-truth labels. While for NA tasks, where answers involve continuous numerical values, we adopt the mean relative accuracy (MRA) metric [66, 43], defined as:
@@ 5.2 Benchmarking MLLM Performance
Table 2 and Table 3 presents the model performance on SpatialBench across five hierarchical levels. Gemini-2.5-pro occupy the top of the ranking and show substantially higher overall scores than any other models. This gap is most pronounced on high-level tasks such as symbolic reasoning, causality, and planning. Nevertheless, several open-source series (e.g., Qwen variants) reach competitive performance on lower and mid levels, indicating that open models can achieve strong perceptual and relational understanding with carefully designed training paradigms. For open-source models, a clear correlation appears between model scale and average performance: larger models generally achieve higher overall scores, showing that scale remains an important factor. However, size alone does not guarantee better performance. Within the same model family, different versions (such as instruction-tuned vs. thinking mode) can show clear performance gaps. This suggests that architectural design and the integration strategy for visual and linguistic information plays a crucial role in spatial cognitive ability.
Figure 5: Comparison of models across spatial cognitive levels.
Fig. 5 illustrates the average performance of several representative models across the five cognitive levels. It can be observed that a cross the board, observation and topology tasks are relatively easier for MLLMs. Most models achieve substantially higher scores on tasks such as object counting, size and distance estimation, and simple topological queries. By contrast, higher-level abilities remain challenging for models, including symbolic reasoning, causality, and visual planning. On these tasks, average performance drops noticeably and variance between models increases. This pattern suggests that while current MLLMs can reliably extract visual evidence and reason about basic relations, they struggle to (a) convert perceptual inputs into robust symbolic rules, (b) infer causal or dynamic consequences accurately, and (c) generate multi-step and convincing plans for a given objective.
@@ 5.3 One-Shot Evaluation
We select several representative models from the one-shot evaluation, including two proprietary models: Gemini-2.5-pro (the best-performing) and GPT-5-chat-latest (the weakest), as well as a strong open-source model Qwen3-VL-235B-A22B-Instruct. Gemini-2.5-pro shows a decline in performance, primarily across the intermediate cognitive levels. In contrast, GPT-5-chat-latest and Qwen3-VL demonstrate substantial improvements. Although these models still trail Gemini in absolute scores, their performance under prompting approaches that of the previously strongest model. This result indicates that even lower-performing models can rapidly enhance specific spatial reasoning abilities when given minimal in-context guidance, highlighting the potential of one-shot prompting to boost MLLMs’ higher-level cognitive capabilities. The results suggest that some powerful models, such as Gemini 2.5 Pro, can reason efficiently and accurately from context on their own, performing better than when given human-guided prompts. This indicates that their built-in contextual understanding is particularly well-suited for spatial intelligence tasks. By contrast, GPT‑5 appears comparatively weaker in intrinsic spatial logic, yet achieves more advanced spatial intelligence by leveraging the GPT series’ steadily improving linguistic capabilities [1]. Similarly, Qwen3-VL, which demonstrates strong text-based reasoning skills [51], benefits from one-shot prompting in a way comparable to GPT‑5, highlighting the role of linguistic reasoning in enhancing spatial cognition under minimal guidance.
@@ 5.4 Benchmarking Human Performance
To quantify the gap between MLLMs and human-level intelligence across cognitive dimensions, we conduct a human benchmark experiment. In this experiment, 33 human participants are presented with SpatialBench, and they provide answers directly without guidance, allowing us to measure their unaided spatial reasoning and planning abilities. The results in Table 2 and Table 3 indicate that human achieve near-perfect performance across nearly all tasks with an overall score of 96.40. Humans perform particularly well on higher-level tasks: symbolic reasoning, causality, and planning, and all achieve essentially 100% accuracy, while even lower-level observation and topology tasks maintain strong performance. Across all levels, humans exhibit not only higher absolute accuracy but also more consistent performance, reflecting robust generalization and contextual understanding that MLLMs have yet to fully achieve. These findings underscore the substantial gap that still exists between human and machine spatial intelligence. While MLLMs show promising abilities in extracting visual information and reasoning about simple topological relations, they remain far from matching human performance in tasks that require integrating high-level reasoning. This test thus provides a clear target for future model development, highlighting the need to enhance multi-step reasoning and context-sensitive planning in MLLMs.
Fig. 6 shows an example of the differences between Gemini and human. The human focuses on the key directional cue: the turning path of the white Volvo, and quickly eliminates irrelevant options based on spatial orientation, demonstrating strong goal-directed and spatially grounded reasoning. In contrast, Gemini describes the entire scene in a more exhaustive but unfocused manner, mentioning many vehicles and areas without identifying the crucial spatial relationship. This suggests that while MLLMs can recognize objects and describe scenes accurately, they often lack selective attention and directional understanding, leading them to infer by association rather than by reasoning about movement and geometry.
5.5 Key Insights from Evaluation
Our evaluation reveals several critical limitations in the spatial cognition of current MLLMs. First, models inherently struggle to reconstruct and maintain continuous spatial scenes from dynamic visual inputs. Poor-performing models frequently misinterpret relative directions or confuse sequential positions, while even advanced models make notable errors in continuous scene tracking and discerning subtle geometric relations. This highlights a fundamental vulnerability in capturing long-range spatial dependencies. Second, our findings empirically validate that spatial cognition in MLLMs is strictly hierarchical and tightly coupled: success in higher-order reasoning is fundamentally bottlenecked by accurate lower-level perception. Performance precipitously declines as cognitive complexity increases, with most models exhibiting cascading failures from the third level onward. To bridge these gaps, future work should leverage high-quality datasets structured with progressive cognitive tasks, incorporate explicit spatial representations, such as scene graphs or 3D geometric priors, and adopt advanced training paradigms, including curriculum learning and agent-based interactive environments, to intrinsically promote multi-step reasoning and long-horizon planning. Furthermore, establishing closed-loop feedback mechanisms, where high-level reasoning and planning errors are utilized to iteratively refine low-level perceptual modules, could significantly mitigate the risk of cascading spatial hallucinations.